OpenStudy (anonymous):

The motion of a particle connected to a spring is described by x = 2 sin (4π t). At what time (in s) is the potential energy equal to the kinetic energy? A.) 0.03 B.) 0.06 C.) 0.08 D.) 0.04 E.) 0.13

OpenStudy (sidsiddhartha):

0.04

OpenStudy (sidsiddhartha):

because sin4pi t=1/2

OpenStudy (praxer):

I think it is 0.06 second cause if the potential energy curve is following a cosine transition than the KE will follow a sine transition or vice verse so the term 4pi t = pi/4 t=1/16=0.0625 Siddhartha you made a wrong in assuming sin4pi t=1/2 it should be 1/root2 http://image.mathcaptain.com/cms/images/41/double-angle-formula-for-cosine.jpg see it once...:) @EducateMe understood.

OpenStudy (praxer):

this is the graphical relation between potential energy and kinetic energy.